The scores of students on the SAT college entrance examinations at a certain high school had a normal distribution with mean μ=553.3 and standard deviation σ=28.6.Round z-scores to 2 decimal places and give probabilities to 4 decimal places.

(a) What is the probability that a single student randomly chosen from all those taking the test scores 558 or higher?
ANSWER:



For parts (b) through (d), consider a simple random sample (SRS) of 30 students who took the test.

(b) What are the mean and standard deviation of the sample mean score x¯, of 30 students?
The mean of the sampling distribution for x¯is:


The standard deviation of the sampling distribution for x¯ is:

(c) What z-score corresponds to the mean score x¯of 558?
ANSWER:


(d) What is the probability that the mean score x¯of these students is 558 or higher?
ANSWER:
Answer #1

Solution! Given that, l-553.3 6-28.6 PL > 558) =1- PCX 2558) =l-p | Xo4 < 558 - 553.3 28.6 -1-e [ z ceny =I-P 4. 28.6 =l- P zDate n=30 4x=553.3 28.6 5.22 6x = 2836 = 5.22 V30 The sample mean ux = 55.3 The standard devition GX = 5.22 L P ( X<558) X-isP(x > 558) 31-PC 5558) =1-2 X-ux <558-553.3 I 6x 5.22 =1-P 4.7 5.22 =1-P za 42 EL-e [za 0.90] -0.8159 = 0.1841 - - P( x2 5S 8